Given Q=4LK +L^2 , Labour cost is 2 and cost is 1, find the maximum out put and the level k and L at which it is achieve when total input cost are fixed at $105





B. Verify that the ratio of marginal product to price is the same for both input at the optimiub

1
Expert's answer
2021-11-07T19:40:03-0500
"\frac {\delta Q}{\delta K}=4L"

"\frac{\delta Q}{\delta L}=4K+2L"

"2L+K=105"

"L=30"

"K=45""\frac{4L}{4K+2L}=\frac{120}{180+60}=0.5"